If the roots of the equation p(q-r)x^2+q(r-p)x+r(p-q)=0 are equal.show that 1/p+1/r=2/q.

Grade:11

1 Answers

Ram Prakash Verma
7 years ago
Take x = 1Root × root = coffeicient of x ÷ coffeicient of x^2r(p-q) = q(r-p)rp+pq= 2prDivide whole by pqr1/p + 1/r = 2/q.........................................